# Large-Scale Galaxy Correlations from the DESI First Data Release

### v1

@@ 1 Introduction
A key question, debated for decades (de Vaucouleurs, 1970; Pietronero, 1987; Davis, 1997; Pietronero et al., 1997; Sylos Labini et al., 1998; Wu et al., 1999; Joyce et al., 2005; Hogg et al., 2005; Sylos Labini et al., 2007; Sarkar et al., 2009; Sylos Labini, 2011; Scrimgeour et al., 2012; Pandey, 2013; Alonso et al., 2014; Whitbourn & Shanks, 2014; Sylos Labini et al., 2014; Conde-Saavedra et al., 2015; Alonso et al., 2015; Sarkar & Pandey, 2016; Pandey & Sarkar, 2015; Shirokov et al., 2016; Laurent et al., 2016; Park et al., 2017; Ntelis et al., 2017; Gonçalves et al., 2018a, b, c; Heinesen, 2020; Teles et al., 2021; De Marzo et al., 2021; Kim et al., 2022; Andrade et al., 2022; Dias et al., 2023; Valade et al., 2024; Pashapour-Ahmadabadi et al., 2025; Einasto, 2025a, b), concerns the scale at which the galaxy distribution can be considered statistically uniform. Determining this scale is essential for assessing whether the observed structures are compatible with current models of galaxy formation. Addressing this issue requires statistical methods that are free of strong assumptions and based on the most conservative hypotheses, in order to minimize luminosity selection biases and finite-size effects. In this respect, one must consider several subtle points.
@@
  • Galaxy counts in magnitude-limited samples from deep, narrow pencil-beam surveys at $z\approx 1$, with limited angular coverage — e.g., the WiggleZ Dark Energy Survey (Scrimgeour et al., 2012);
@@ 3.1 Sample geometry and boundary conditions
3.2 Bright Galaxy Sample (BGS)
@@ 4.2 The average conditional density in the Cold Dark Matter model
The normalization constant $A$ is chosen such that $\overline{\xi}(r_{0})=1$ for $r_{0}=9$ and $14$ Mpc/h/h, respectively, where $\xi(r)$ is the correlation function — the Fourier conjugate of $P(k)$ — and
@@ 5 Conclusions
Acknowledgements
This work is dedicated to the memory of Yurij V. Baryshev.
F.S.L. is grateful to Michalel Joyce for a list of comments and suggestions and to Giordano De Marzo and Andrea Gabrielli for many insightful and stimulating discussions.
This research used data obtained with the Dark Energy Spectroscopic Instrument (DESI). DESI construction and operations is managed by the Lawrence Berkeley National Laboratory. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of High-Energy Physics, under Contract No. DE–AC02–05CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSF’s National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Humanities, Science and Technology of Mexico (CONAHCYT); the Ministry of Science and Innovation of Spain (MICINN), and by the DESI Member Institutions: www.desi.lbl.gov/collaborating-institutions. The DESI collaboration is honored to be permitted to conduct scientific research on I’oligam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U.S. National Science Foundation, the U.S. Department of Energy, or any of the listed funding agencies.

### v2

@@ 1 Introduction
A key question, debated for decades (de Vaucouleurs, 1970; Pietronero, 1987; Lemson & Sanders, 1991; Coleman & Pietronero, 1992; Davis, 1997; Pietronero et al., 1997; Sylos Labini et al., 1998; Wu et al., 1999; Joyce et al., 2005; Hogg et al., 2005; Sylos Labini et al., 2007; Sarkar et al., 2009; Sylos Labini, 2011; Scrimgeour et al., 2012; Pandey, 2013; Alonso et al., 2014; Whitbourn & Shanks, 2014; Sylos Labini et al., 2014; Conde-Saavedra et al., 2015; Alonso et al., 2015; Sarkar & Pandey, 2016; Pandey & Sarkar, 2015; Shirokov et al., 2016; Laurent et al., 2016; Park et al., 2017; Ntelis et al., 2017; Gonçalves et al., 2018a, b; Heinesen, 2020; Teles et al., 2021; De Marzo et al., 2021; Kim et al., 2022; Andrade et al., 2022; Dias et al., 2023; Valade et al., 2024; Pashapour-Ahmadabadi et al., 2025; Einasto, 2025a, b), concerns the scale at which the galaxy distribution can be considered statistically uniform. Determining this scale is essential for assessing whether the observed structures are compatible with current models of galaxy formation. Addressing this issue requires statistical methods that are free of strong assumptions and based on the most conservative hypotheses, in order to minimize luminosity selection biases and finite-size effects. In this respect, one must consider several subtle points.
@@
  • Galaxy counts in magnitude-limited samples from deep, narrow pencil-beam surveys at $z\approx 1$, with limited angular coverage — e.g. the WiggleZ Dark Energy Survey (Scrimgeour et al., 2012). We note that the procedure adopted in the WiggleZ analysis, which corrects counts-in-spheres in a non–volume-limited sample by normalizing to the expectation for a homogeneous distribution with the same completeness, implicitly assumes large-scale homogeneity. As a result, the method effectively builds in the conclusion it seeks to demonstrate, amounting to a form of circular reasoning. This correction scheme has subsequently been adopted in later studies (e.g. Gonçalves et al., 2021);
@@ 3.1 Sample geometry and boundary conditions
3.2 Bright Galaxy Sample
@@ 4.2 The average conditional density in the Cold Dark Matter model
The normalization constant $A$ is chosen such that $\overline{\xi}(r_{0})=1$ for $r_{0}=9$ and $14$ Mpc/h/h, respectively, where $\xi(r)$ is the correlation function—the Fourier conjugate of P(k)P(k)—and
@@ 5 Conclusions
Data Availability
The DESI data employed in this work are publicly available through the DESI collaboration. The analysis pipeline developed for this study is available from the authors upon request.
Acknowledgements.
This work is dedicated to the memory of Yurij V. Baryshev. F.S.L. is grateful to Michalel Joyce for a list of comments and suggestions and to Giordano De Marzo and Andrea Gabrielli for many insightful and stimulating discussions. This research used data obtained with the Dark Energy Spectroscopic Instrument (DESI). DESI construction and operations is managed by the Lawrence Berkeley National Laboratory. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of High-Energy Physics, under Contract No. DE–C02–05CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSF’s National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Humanities, Science and Technology of Mexico (CONAHCYT); the Ministry of Science and Innovation of Spain (MICINN), and by the DESI Member Institutions: www.desi.lbl.gov/collaborating-institutions. The DESI collaboration is honored to be permitted to conduct scientific research on I’oligam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U.S. National Science Foundation, the U.S. Department of Energy, or any of the listed funding agencies.